OpenStudy (anonymous):

1.2258 = 9.8(sin θ) I do not know how to solve trigonometric equations and would really appreciate any help!:) I need to solve for theta.

OpenStudy (anonymous):

\[\theta =\frac{180 \sin ^{-1}\left(\frac{1.2258}{9.8}\right)}{\pi }=7.18547 {}^{ \ \circ} \]